### v1

@@ 2.3 Human Pose as HOI Cues
Recognizing the importance of human pose in understanding human behavior and intention, researchers have explored various methods [11, 34, 35, 36, 37, 38, 57, 58] to extract and leverage pose features to enhance interaction representations. For example, Park et al. [11] designed a pose-conditioned graph neural network that utilizes local features of human joints to update human node encoding to contain richer local information. Qiao et al. [34] focused on extracting geometric features, such as human posture and object position, to supplement visual features to improve robustness in partially occluded scenes. Li et al. [35] emphasized the unique characteristics of human body parts related to interactivity and proposed a hierarchical learning framework based on instance-level and part-level body features. However, these works mainly focus on the local pose features of the target person, neglecting global clues from other people in the image. To overcome this limitation, Wu et al. [58] introduced body-part saliency maps to capture multi-person features and learn the overall relationship between different body parts. Nevertheless, most of these methods rely on off-the-shelf human pose estimators or body part detectors, significantly increasing complexity and computational costs. Moreover, these models are typically trained on third-person datasets, making their application challenging in first-person scenarios. To address these issues, our work leverages the geometric robustness of global hand pose features to provide crucial complementary information to visual features, deepening our understanding of the complex dynamics of Ego-HOI under partial occlusion. We integrate hand pose estimation into our Ego-HOI detection pipeline, sharing weights with the hand detection branch. This integration not only addresses generalization limitations but also reduces the computational burden, making the entire system more efficient and practical.
@@ 3 Our Ego-HOIBench Benchmark
Ego-HOIBench is an egocentric image dataset explicitly annotated for Ego-HOI detection research. The dataset provides high-quality ground truth annotations for hand-object pair detection and interaction recognition across all frames. Hand and object annotations contain multiple $\left({class,bbox}\right)$ tuples, where $class$ indicates the hand side (left or right) or object category, and $bbox$ denotes a bounding box determined by the coordinates of its top-left and bottom-right corners. Interaction annotations specify the exact action category performed by each hand-object pair. Combined with original hand pose annotations, our Ego-HOIBench dataset provides rich details for studying human-object interactions in egocentric vision.
@@ 3.1 Generation Steps
We perform the following steps to acquire images and generate annotations for our Ego-HOIBench benchmark. Given an untrimmed RGB-D video sequence derived from the HOI4D dataset [30], we begin by extracting the intermediate frames from each action clip, based on the identified start and end timestamps, as these frames effectively capture sufficient information. Then, the intermediate frame’s mask regions are associated with the corresponding object categories. According to the definition of the Ego-HOI detection task, we focus only on the active objects in current frames. By analyzing the task information, we restrict the possible categories of active objects and filter out irrelevant objects. To avoid meaningless component segmentation, we merge different components of the same objects, e.g., the scissors’ left and right parts, and the safe’s body and door. Subsequently, we convert the mask regions into bounding boxes. A hand-object pair’s bounding boxes and categories are combined with the corresponding action category to form a complete Ego-HOI triplet annotation, i.e., <human hand, verb, object>. Since subtle errors in pixel-level masks can lead to considerable deviations in corresponding bounding boxes, we employ human experts to double-check and ensure accurate annotations. The label correction work is time-consuming, requiring a combined effort of twenty person-days. The entire dataset generation process spans approximately one and a half months. The extracted intermediate frames and their annotations constitute the Ego-HOIBench dataset. The Ego-HOIBench dataset is further divided into training and test sets. We split the frames according to their video identities to ensure no overlap of object entities in the training and test sets. With a split ratio of 80%:20%, we finally obtain 22,088 training frames and 5,487 testing frames.
@@ 3.2 Dataset Statistics
Figure 2: Distributions of objects (top) and verbs (bottom), sorted by instance count. Some categories appear significantly more frequently than others.
Among the observed instances, the vast majority (91.4%) rely on the right hand alone, while fewer cases (8.2%) use both hands. Even fewer cases operate with the left hand alone, accounting for only 0.4%. Fig. 2 shows the distributions of object and verb categories under the instance-level setting, the definition of which will be illustrated in Sec. 3.3. The instance number of different object categories show a significant span, ranging from 2630 to 8. This wide distribution range is also reflected in the number of verbs. The triplet combination of hands, verbs, and objects further exacerbates the data imbalance, reflecting the natural distribution of HOIs in the real world. This characteristic makes Ego-HOIBench a distinctive and challenging benchmark for Ego-HOI detection, presenting challenges that are closely related to practical applications.
Fig. LABEL:fig3 shows the co-occurrence between verbs and objects (e.g., Open and Drawer). Each square in the heat map directly reflects the number of instances involving a particular verb-object pair, with darker colors indicating more corresponding instances. Our dataset contains various distinctive co-occurrence scenes, where some specific verbs, such as Turn on only associated with the object Lamp, and Press only co-occurs with Trash can. The co-occurrence between objects (including hands) and verbs highlights the feasibility of using this information to suppress negative prediction candidates. This suppression scheme closely mirrors human decision-making processes and is therefore frequently employed during the model inference stage [11, 43].
Table 2: Statistics of instances and Ego-HOI triplet categories at different occlusion ratios in our Ego-HOIBench Dataset.
#Instances #Triplet Categories
Split #Frames #Total Ins. @0$\sim$0.2 @0.2$\sim$0.4 @0.4$\sim$0.6 @0.6$\sim$0.8 @0.8$\sim$1 @0$\sim$0.2 @0.2$\sim$0.4 @0.4$\sim$0.6 @0.6$\sim$0.8 @0.8$\sim$1
@@
Table 2 provides detailed statistics on the number of instances and Ego-HOI triplet categories at various occlusion ratios for the training and test sets in our Ego-HOIBench dataset. The occlusion ratio of an Ego-HOI instance is calculated by dividing the area of the object occluded by hands and other objects by the area of its bounding box. In our dataset, occlusion is common, with roughly half of the instances having at least 20% of their area occluded and about 20% having an occlusion ratio over 40%. The high occlusion ratios increase the difficulty of detection and recognition and affect the model’s generalization and robustness. In addition, the number of triplet categories decreases significantly as occlusion increases. This phenomenon is closely related to the physical size of the objects. Larger objects, such as cabinets and chairs, are typically only slightly obscured by hands or other objects. In contrast, smaller objects, like staplers and bowls, are more prone to varying degrees of occlusion.
@@ 3.3 Ego-HOI Detection Tasks
Following a third-person perspective, HOI is defined as a triplet containing a person, a verb, and an object [14, 59]. It assumes a one-to-one correspondence between these three elements. Since people, as interacting subjects, remain constant, most HOI detection models disregard subject identification but focus solely on the localization of humans. In the context of egocentric vision, Ragusa et al. [25] described interactions with multiple objects using <verb, objects>, completely ignoring the role of the human hand in the interaction. However, a person’s left and right hands can independently perform different interactions or collaborate on a single interaction, making it an oversimplification to treat a person’s hands as a single, unchanging entity. Furthermore, human hands are not passive in activities but actively influence and shape the nature of interactions. Therefore, a comprehensive understanding of hand factors is indispensable for Ego-HOI detection.
Building upon the aforementioned understanding, we redefine Ego-HOI as the <hands, verb, objects> triplet. We emphasize the significance of hands in interactions and consider comprehending their categories and positions as essential for understanding egocentric interaction dynamics. In light of the definition of Ego-HOI, we present two detection tasks to evaluate the model’s capacity to comprehend interactions at the instance level and the abstract image level.
Figure 4: Examples of general, instance-level specific, and image-level specific hand-object interactions. The instance-level setting focuses on the interaction behaviors involving a single active object, while the image-level setting interprets the interaction holistically from the perspective of the entire image. In the instance-level setting (second row), an image may be parsed into two separate interaction instances: right-hand bind paper and right-hand bind with stapler. In contrast, the image-level setting (last row) defines the image as a unified interaction: right-hand bind paper with stapler. For clarity, the hand category is omitted from the image captions.
3.3.1 Instance-level Ego-HOI Detection Task
Let =$\mathcal{H}=\left\{h_{r},h_{l}\right\}$, $\mathcal{V}=\left\{v_{1},v_{2},\dots,v_{m}\right\}$, and $\mathcal{O}=\left\{o_{1},o_{2},\dots,o_{n}\right\}$ denote the sets of hands, verbs, and objects, respectively, where $m$ and $n$ are the number of verbs’ and objects’ categories, respectively. We define the prediction target for each instance as follows:
@@ 3.3.1 Instance-level Ego-HOI Detection Task
where $\left(\overline{h_{r}},\overline{h_{l}}\right)$ are the hands engaged in the interaction. There are three situations to consider: right hand only $\left(h_{r},\cdot\right)$, left hand only $\left(\cdot,h_{l}\right)$, and both hands $\left(h_{r},h_{l}\right)$. $v_{i}\subset\mathcal{V}$ is the verb that describes interaction, and $o_{j}\subset\mathcal{O}$ is the object of interest. The annotations for each instance comprise the class labels for the hand(s) and active object, along with their respective bounding boxes and the verb class label. In total, we define 123 Ego-HOI triplet categories consisting of one or two hands, a verb, and an object, e.g., right-hand cut apple and left and right hands dump bucket.
3.3.2 Image-level Ego-HOI Detection Task
The objective of image-level Ego-HOI detection is to deduce the primary interaction within each frame and identify all hands and active objects participating in it. Compared with the instance-level setting, the image-level setting comprehensively considers the objects directly and indirectly involved in the interaction when analyzing triplets. For example, in the instance-level setting, an image may be parsed into two separate interaction instances right-hand bind paper and right-hand bind with stapler. In contrast, the image-level setting defines the image as a unified interaction right-hand bind paper with stapler. This image-level perspective examines and explains interaction behaviors from a broader perspective. To explain the instance-level and image-level Ego-HOI detection tasks more clearly, we show examples of hand-object interactions in Fig. 4.
@@ 4.1 HGIR Architecture
Our HGIR scheme is straightforward yet robust and can be easily integrated with various baseline HOI detection methods, yielding appealing results in the Ego-HOI detection task. The overall architecture of our method is shown in Fig. 5.
@@
Given an input RGB image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, we employ the original baseline HOI detection method to obtain the hand features $\mathbf{H}\in\mathbb{R}^{N\times d}$, the object features $\mathbf{O}\in\mathbb{R}^{N\times d}$, and the interaction features $\mathbf{I}\in\mathbb{R}^{N\times d}$, denoted as $\left(\mathbf{H},\mathbf{O},\mathbf{I}\right)=Baseline\left(\mathbf{X}\right)$. The baseline method can adopt either a unified or decoupled prediction strategy as long as it provides the necessary interaction (i.e., verb) and hand (i.e., subject) representations. Multiple hand pose candidates $\hat{\mathcal{G}}\in\mathbb{R}^{N\times 2N_{g}}$ are estimated based on $\mathbf{H}$, where $N_{g}$ is the number of hand joints. Then, a selection strategy is designed to generate left-hand and right-hand pose proposals, and their geometric features $\mathbf{f}\in\mathbb{R}^{2KN_{g}\left(N_{g}-1\right)}$ are extracted to describe the details of hand structure. Simultaneously, the interactivity refinement block uses the attention mechanism to direct the interaction features focus toward the regional information derived from pose offset prompts $\mathbf{H}^{\mathrm{off}}$. These two features are fused to obtain the ultimate interaction embedding $\mathbf{E}\in\mathbb{R}^{N\times d}$ for classification. Overall, our HGIR scheme exploits the synergy of complementary geometric features and refined interaction features to enhance the ability to perceive interaction dynamics.
@@ 4.2 Hand Pose Estimation
[…]
@@
Combining reference points and offsets instead of directly predicting joint positions offers two key advantages. Firstly, it avoids the complexity of directly finding joint points from the entire image. Secondly, the pose offset-specific features can act as valuable prompts in the subsequent interactivity refinement.
@@ 4.3 Hand Geometry Extraction
$\begin{matrix}\Omega_{l}^{\ast}=\Omega_{l}^{{}^{\prime}}\cup\left\{\mathbf{\hat{g}}_{i}\in\Omega_{r}^{{}^{\prime}}\mid i=1,\dots,K-\left|\Omega_{l}^{{}^{\prime}}\right|\right\}\\ \Omega_{r}^{\ast}=\Omega_{r}^{{}^{\prime}}\cup\left\{\mathbf{\hat{g}}_{i}\in\Omega_{l}^{{}^{\prime}}\mid i=1,\dots,K-\left|\Omega_{r}^{{}^{\prime}}\right|\right\}\end{matrix}$ (5)
[…]
@@ 4.4 Interactivity Refinement
The refiner contains a self-attention layer [50] that focuses on capturing and modeling the intrinsic correlations within the interaction features, obtaining the advanced interaction features ${\mathbf{I}}^{\prime}=\left\{{\mathbf{I}_{i}}^{\prime}\right\}^{N}_{i=1}$. Next, we introduce the pose offset-specific features $\mathbf{H}^{\mathrm{off}}$ as pose prompts to inject pose awareness into the advanced interaction features. Specifically, we feed $\mathbf{H}^{\mathrm{off}}$ into the attention mechanism as keys and values, while ${\mathbf{I}}^{\prime}$ serves as queries. Each output element ${\mathbf{I}_{i}}^{\prime\prime}$ is computed by aggregating all values weighted with attention: ${\mathbf{I}_{i}}^{\prime\prime}={\textstyle\sum_{j}}\alpha_{ij}\left(\mathbf{W}_{v}\mathbf{h}_{j}^{\mathrm{off}}\right)$, where $\alpha_{ij}$ is the normalized attention weight, as follows:

### v2

@@ 2.3 Human Pose as HOI Cues
Human pose information is a powerful cue for understanding human-object interactions, as it encodes both the intent and the fine-grained dynamics of actions. Numerous studies (Gupta et al., 2019; Wan et al., 2019; Zheng et al., 2020; Wu et al., 2022; Qiao et al., 2022; Li et al., 2022; Park et al., 2023) have demonstrated that incorporating pose features can significantly enhance interaction recognition, especially in challenging scenarios with occlusions or ambiguous visual contexts. For instance, Park et al. (2023) proposed a pose-conditioned graph neural network that enriches human node representations with local joint features, leading to more discriminative interaction modeling. Similarly, Qiao et al. (2022) showed that geometric features, such as body posture and object position, can robustly supplement visual features, improving performance under partial occlusion. Hierarchical frameworks like (Li et al., 2022) further exploit both instance-level and part-level body cues to capture the unique characteristics of interactive body parts. While pose cues are valuable, most existing methods rely on external pose estimators trained on third-person data, which increases complexity and often fails under egocentric occlusion. To address these issues, we embed hand pose estimation directly into the Ego-HOI detection framework, sharing features with the hand detection branch. This unified, end-to-end strategy yields robust, efficient hand geometry cues tailored for egocentric scenarios, improving both accuracy and practicality.
@@ 3 Our Ego-HOIBench Benchmark
Ego-HOIBench is an egocentric image dataset designed specifically for Ego-HOI detection research. It provides high-quality, frame-level annotations for both hand-object pair detection and interaction recognition. Each hand and object is annotated with a tuple $\left({class,bbox}\right)$, where classclass indicates the hand side (left or right) or the object category, and bboxbbox specifies the bounding box using the coordinates of the top-left and bottom-right corners. For every hand-object pair, the dataset includes the corresponding interaction category. In addition, original hand pose annotations are provided, offering detailed information for analyzing human-object interactions from an egocentric perspective.
@@ 3.1 Generation Steps
We generate the Ego-HOIBench dataset through the following steps: given an untrimmed RGB-D video sequence derived from the HOI4D dataset (Liu et al., 2022), we first extract intermediate frames from each action clip, using the annotated start and end timestamps, to ensure that the selected frames capture the key moments of each interaction. Next, we associate the mask regions in each intermediate frame with their corresponding object categories. According to the definition of the Ego-HOI detection task, we focus only on the active objects present in each frame. By analyzing the task information, we restrict the possible active object categories and filter out irrelevant objects. To avoid redundant segmentation, we merge different components of the same object (e.g., the left and right blades of scissors, or the body and door of a safe) into a single object instance. Subsequently, we convert the mask regions into bounding boxes. For each hand-object pair, we combine the bounding boxes and category labels with the corresponding action category to form a complete Ego-HOI triplet annotation, i.e., <human hand, verb, object>. To ensure annotation quality, all automatically generated annotations are manually reviewed and corrected as needed. Independent annotators cross-check the results to guarantee high-quality ground truth. This manual correction process is time-consuming, requiring a total of about twenty person-days. The entire dataset generation process takes approximately one and a half months. The extracted intermediate frames and their annotations constitute the Ego-HOIBench dataset. We further divide the dataset into training and test sets by splitting frames according to their video identities, ensuring that object entities do not overlap between the two sets. Using an 80%/20% split, we obtain 22,088 training frames and 5,487 test frames.
@@ 3.2 Dataset Statistics
Figure 2: Distribution of object (top) and verb (bottom) categories in Ego-HOIBench, sorted by instance count. Most instances are concentrated in a few common categories, while many categories have relatively few examples.
Among all annotated instances, the majority (91.4%) involve only the right hand, while 8.2% involve both hands, and just 0.4% involve only the left hand. Fig. 2 presents the distributions of object and verb categories (defined in Sec. 3.3). The instance number of different object categories varies widely, from as many as 2,630 to as few as 8, and a similar imbalance is observed across verb categories. When considering the full triplet combinations of hands, verbs, and objects, this imbalance becomes even more pronounced, reflecting the natural distribution of HOIs in real-world scenarios. As a result, Ego-HOIBench poses unique and realistic challenges for Ego-HOI detection, making it a valuable and demanding benchmark for practical applications.
Fig. 3 presents a heatmap of verb-object co-occurrence in Ego-HOIBench. Each cell represents the number of instances for a specific verb-object pair, with darker colors indicating higher frequencies. The dataset exhibits a variety of distinctive co-occurrence patterns. These co-occurrence statistics reveal that certain verb-object pairs are much more likely to occur than others. Leveraging this information can help suppress unlikely or negative prediction candidates during model inference, a strategy that aligns with human reasoning and has been widely adopted in recent HOI detection methods. (Gao et al., 2020; Park et al., 2023).
Figure 3: Heatmap of verb-object co-occurrence in Ego-HOIBench. Each cell indicates the number of instances for a specific verb-object pair; darker colors represent higher frequencies.
Table 2: Number of instances and Ego-HOI triplet categories in the Ego-HOIBench dataset, grouped by object occlusion ratio. The occlusion ratio is defined as the proportion of the object area occluded by hands or other objects.
#Instances by Occlusion Ratio #Triplet Categories by Occlusion Ratio
[…]
@@
Table 2 provides a detailed breakdown of Ego-HOI instance counts and unique triplet categories in the Ego-HOIBench training and test sets, grouped by object occlusion ratio. Here, the occlusion ratio is defined as the fraction of an object’s bounding box area that is occluded by hands or other objects. Occlusion is common in our dataset: about half of all instances have at least 20% occlusion, and roughly 20% of instances are more than 40% occluded. Higher occlusion ratios make both detection and recognition more difficult, posing significant challenges for model robustness and generalization. Additionally, as occlusion increases, the number of unique triplet categories drops sharply. This is mainly because large objects (such as cabinets or chairs) are usually partially occluded, while small objects (such as staplers or bowls) are more likely to suffer from different degrees of occlusion.
@@ 3.3 Ego-HOI Detection Tasks
Unlike the third-person HOI detection, where the subject is typically treated as a generic “person” and subject recognition is ignored (Ning et al., 2023; Dong et al., 2022), egocentric vision fundamentally requires explicit modeling of the hands. In first-person scenarios, the left and right hands often act independently or collaboratively, each contributing to the interaction. Existing egocentric benchmarks, such as MECCANO (Ragusa et al., 2023), overlook subjects and annotate only verbs and objects, completely omitting explicit hand categories and locations. However, in egocentric HOI, the identity (left/right) and precise position of the hands are critical for correctly interpreting the interaction, as they directly determine the semantics of the action. Therefore, we define Ego-HOI as a <hand(s), verb, object> triplet, explicitly placing hands at the center of egocentric interaction understanding. This formulation is essential for accurate Ego-HOI detection and enables more fine-grained and robust modeling of real-world egocentric activities.
Figure 4: Illustration of <hand(s), verb, object> triplet annotations for Ego-HOI detection. Each interaction instance is annotated with bounding boxes and categories for the hand(s) and active object, as well as the verb class label.
Let $\mathcal{H}=\{h_{r},h_{l}\}$ denote the set of hands (right and left), $\mathcal{V}=\{v_{1},v_{2},\dots,v_{m}\}$ the set of verbs, and $\mathcal{O}=\{o_{1},o_{2},\dots,o_{n}\}$ the set of objects, where $m$ and $n$ are the numbers of verb and object categories, respectively. For each interaction instance, the prediction target is defined as:
@@
where $\left(\overline{h_{r}},\overline{h_{l}}\right)$ indicates which hand(s) are involved in the interaction. There are three possible cases: right hand only $\left(h_{r},\cdot\right)$, left hand only $\left(\cdot,h_{l}\right)$, or both hands $\left(h_{r},h_{l}\right)$. $v_{i}\in\mathcal{V}$ is the verb describing the interaction, and $o_{j}\in\mathcal{O}$ is the object being interacted with. Each instance is annotated with the class labels and bounding boxes for the hand(s) and active object, as well as the verb class label, as illustrated in Fig. 4. In total, we define 123 unique Ego-HOI triplet categories, each consisting of one or two hands, a verb, and an object (e.g., right-hand cut apple, left and right hands dump bucket).
@@ 4.1 HGIR Architecture
Our HGIR scheme is straightforward yet robust and can be easily integrated with various baseline HOI detection methods, yielding appealing results in the Ego-HOI detection task. The overall architecture of our method is shown in Fig. 5. Given an input RGB image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, we employ the original baseline HOI detection method to obtain the hand features $\mathbf{H}\in\mathbb{R}^{N\times d}$, the object features $\mathbf{O}\in\mathbb{R}^{N\times d}$, and the interaction features $\mathbf{I}\in\mathbb{R}^{N\times d}$, denoted as $\left(\mathbf{H},\mathbf{O},\mathbf{I}\right)=Baseline\left(\mathbf{X}\right)$. The baseline method can adopt either a unified or decoupled prediction strategy as long as it provides the necessary interaction (i.e., verb) and hand (i.e., subject) representations. Multiple hand pose candidates $\hat{\mathcal{G}}\in\mathbb{R}^{N\times 2N_{g}}$ are estimated based on $\mathbf{H}$, where $N_{g}$ is the number of hand joints. Then, a selection strategy is designed to generate left-hand and right-hand pose proposals, and their geometric features $\mathbf{f}\in\mathbb{R}^{2KN_{g}\left(N_{g}-1\right)}$ are extracted to describe the details of hand structure. Simultaneously, the interactivity refinement block uses the attention mechanism to direct the interaction features focus toward the regional information derived from pose offset prompts $\mathbf{H}^{\mathrm{off}}$. These two features are fused to obtain the ultimate interaction embedding $\mathbf{E}\in\mathbb{R}^{N\times d}$ for classification. Overall, our HGIR scheme exploits the synergy of complementary geometric features and refined interaction features to enhance the ability to perceive interaction dynamics.
@@ 4.2 Hand Pose Estimation
where $\delta$ and $\sigma$ are the sigmoid and softmax operations, respectively. $\left|\mathcal{H}\right|$ denotes the number of hand category set, and the additional class represents the background class (no object). The predicted category $\hat{c}_{i}^{\mathrm{h}}$ and score $\hat{s}_{i}^{\mathrm{h}}$ are given by $\underset{k}{\arg\max\,}\mathbf{\hat{p}}_{i,k}^{\mathrm{h}}$ and $\underset{k}{\max\,}\mathbf{\hat{p}}_{i,k}^{\mathrm{h}}$.
@@
Combining reference points and offsets, rather than directly regressing joint positions, provides two significant advantages. First, it constrains the search space for joint localization to the vicinity of detected hands, which greatly improves accuracy and efficiency compared to unconstrained regression over the entire image. Second, the explicit modeling of pose offsets enables the network to learn fine-grained, context-dependent hand configurations, which can be leveraged as informative prompts in the subsequent interactivity refinement module. This strategy not only enhances the robustness of hand pose estimation under occlusion but also facilitates more precise and discriminative interaction reasoning.
@@ 4.3 Hand Geometry Extraction
$\Omega_{l}^{\ast}=\Omega_{l}^{{}^{\prime}}\cup\left\{\mathbf{\hat{g}}_{i}\in\Omega_{r}^{{}^{\prime}}\mid i=1,\dots,K-\left|\Omega_{l}^{{}^{\prime}}\right|\right\}$ (5)
$\Omega_{r}^{\ast}=\Omega_{r}^{{}^{\prime}}\cup\left\{\mathbf{\hat{g}}_{i}\in\Omega_{l}^{{}^{\prime}}\mid i=1,\dots,K-\left|\Omega_{r}^{{}^{\prime}}\right|\right\}$ (6)
where $\left|\cdot\right|$ indicates ther number of the set. In this manner, regardless of the number of valid candidates, the final sets $\Omega_{l}^{\ast}$ and $\Omega_{r}^{\ast}$ will each contain exactly $K$ proposals.
@@ 4.4 Interactivity Refinement
The refiner contains a self-attention layer (Carion et al., 2020) that focuses on capturing and modeling the intrinsic correlations within the interaction features, obtaining the advanced interaction features ${\mathbf{I}}^{\prime}=\left\{{\mathbf{I}_{i}}^{\prime}\right\}^{N}_{i=1}$. Next, we introduce the pose offset-specific features $\mathbf{H}^{\mathrm{off}}$ as pose prompts to inject pose awareness into the advanced interaction features. Specifically, we feed $\mathbf{H}^{\mathrm{off}}$ into the attention mechanism as keys and values, while $\mathbf{I}$$^{\prime}$ serves as queries. Each output element $\mathbf{I}_{i}$$^{\prime\prime}$ is computed by aggregating all values weighted with attention: ${\mathbf{I}_{i}}^{\prime\prime}={\textstyle\sum_{j}}\alpha_{ij}\left(\mathbf{W}_{v}\mathbf{h}_{j}^{\mathrm{off}}\right)$, where $\alpha_{ij}$ is the normalized attention weight, as follows: